# The ELM Neuron: an Efficient and Expressive

### v1

@@ Appendix C Additional Results
Figure S1: The ELM neuron is a computationally efficient model of cortical neuron. Similar figure to 2c and 2d, however, displaying FLOPs required to do inference on a single sample. a) and b) Voltage and spike prediction performance of the respective surrogate models. While previous works required around 10M parameters to make accurate spike predictions using a TCN Beniaguev et al. [2021], an LSTM baseline is able to do it with 266K, our ELM neuron model requires merely 53K, and our (branch) ELM neuron only a humble 8K, simultaneously achieving much better voltage prediction performance than the TCN. A throughput optimized ELM neuron implementation can potentially reduce the required FLOPs even further.
@@
Figure S2: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an (branch) ELM neuron with AUC $\approx 0.992$, with default hyperparameters. The number of removed divergent runs marked with $1^{*}$. a) We find above memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich [2004], Dayan and Abbott [2005]. b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart and Spruston [2015], Jones and Kording [2022], Larkum [2022]. c) Allowing greater updates to the memory units is beneficial, however, too large ones increase training instability. d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with the default range being 1ms-150ms. Timescales around 25ms-50ms seem to be the most useful (matching the typical membrane timescale in the cortex Dayan and Abbott [2005]); however, a lack can be partially compensated by longer timescales, even better than by the vanilla ELM. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$.
@@
Figure S3: The ELM neuron performs well on typical spiking datasets. The following results are on the original Spiking Heidelberg Digits dataset Cramer et al. [2020b]. a-c) The ELM neuron consistently outperforms a classic LSTM, especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. Our Leaky Integrate and Fire (LIF) Spiking Neural Network (SNN), however, does manage to achieve decent performance for intermediate bin sizes around 20, in contrast to the SHD-Adding dataset (see Figure 5).
@@ Appendix D Additional Visualizations
Figure S4: Visualization of ELM neuron dynamics. Extended visualization of Figure 2b for an ELM neuron achieving around 0.992 AUC. a) The synaptic input as the neuron model receives it, with excitatory input ($+1$) marked in red, and inhibitory input ($-1$) marked in blue. b) The ELM neurons predictions and the ground-truth targets, for a particularly challenging example. c) A random subset of 20 synapses are visualized. Synapses receiving negative input will be deflected downwards. d) All 20 memory values are visualized. Some fluctuate more rapidly than others, typically proportional to their memory timescales.
Appendix E Reduced Implementation In Pytorch
import math
import torch
import torch.nn as nn
# NOTE: the 1e-6 used throughout are to avoid numerical instabilities
# NOTE: the tau_m bounds are enforced using a scaled sigmoid
# NOTE: the w_s bounds are enforced using relu (originally exp)
def scaled_sigmoid(x, lower_bound, upper_bound):
    return (upper_bound - lower_bound) * torch.sigmoid(x) + lower_bound
def inverse_scaled_sigmoid(x, lower_bound, upper_bound):
    x = torch.clamp(x, lower_bound + 1e-6, upper_bound - 1e-6)
    return torch.log((x - lower_bound) / (upper_bound - x))
def custom_tanh(x):
    return torch.tanh(x * 2 / 3) * 1.7159
class MLP(nn.Module):
    """Multi-layerperceptron(MLP)model."""
    def __init__(self, input_size, hidden_size, num_output, num_layers):
        super(MLP, self).__init__()
        next_input_size = input_size
        layers = []
        for i in range(num_layers):
            layers.append(nn.Linear(next_input_size, hidden_size))
            layers.append(nn.ReLU())
            next_input_size = hidden_size
        layers.append(nn.Linear(next_input_size, num_output))
        self.network = nn.Sequential(*layers)
    def forward(self, x):
        return self.network(x)
class ELM(nn.Module):
    """ExpressiveLeakyMemory(ELM)neuronmodel."""
    def __init__(self, num_synapse, num_output, num_memory=100,
                 mlp_num_layers=1, memory_tau_min=10, memory_tau_max=1000,
                 lambda_value=3.0, tau_s_value=5.0, w_s_value=0.5, delta_t=1.0,
                 learn_memory_tau=False,mlp_hidden_size=None,use_num_branch=None):
        super(ELM, self).__init__()
        self.delta_t = delta_t
        self.lambda_value = lambda_value
        self.learn_memory_tau = learn_memory_tau
        self.memory_tau_min = memory_tau_min
        self.memory_tau_max = memory_tau_max
        # configuration in case of using branches
        self.use_num_branch = use_num_branch
        if use_num_branch is not None:
            assert num_synapse % use_num_branch == 0, "mustbeexactlydivisible"
            num_mlp_input = use_num_branch + num_memory
        else:
            num_mlp_input = num_synapse + num_memory
        # initialization of model parameters
        mlp_hidden_size = mlp_hidden_size if mlp_hidden_size else 2 * num_memory
        self.mlp = MLP(num_mlp_input, mlp_hidden_size, num_memory, mlp_num_layers)
        self.w_s = nn.parameter.Parameter(torch.full((num_synapse,), w_s_value))
        self.w_y = nn.Linear(num_memory, num_output)
        # initialization of time constants and decay factors
        self.tau_s = torch.full((num_synapse,), tau_s_value)
        self.kappa_s = torch.exp(-delta_t / torch.clamp(self.tau_s, min=1e-6))
        self._proto_tau_m = torch.logspace(
            math.log10(memory_tau_min + 1e-6),
            math.log10(memory_tau_max - 1e-6),
            num_memory,
        self._proto_kappa_m = torch.exp(
            -delta_t / torch.clamp(self._proto_tau_m, min=1e-6)
        # configuration in case memory tau is learnable
        if self.learn_memory_tau:
            self._proto_tau_m = inverse_scaled_sigmoid(
                self._proto_tau_m, memory_tau_min, memory_tau_max
            self._proto_tau_m = nn.parameter.Parameter(self._proto_tau_m)
    @property
    def tau_m(self):
        # ensure that tau_m is always up to date
        if self.learn_memory_tau:
            return scaled_sigmoid(
                self._proto_tau_m, self.memory_tau_min, self.memory_tau_max
        else:
            return self._proto_tau_m
    @property
    def kappa_m(self):
        # ensure that kappa_m is always up to date
        if self.learn_memory_tau:
            return torch.exp(-self.delta_t / torch.clamp(self.tau_m, min=1e-6))
        else:
            return self._proto_kappa_m
    def dynamics(self, x, s_prev, m_prev, kappa_m):
        # compute the dynamics for a single timestep
        batch_size = x.shape[0]
        s_t = self.kappa_s * s_prev + torch.relu(self.w_s) * x
        if self.use_num_branch is not None:
            syn_input = s_t.view(batch_size, self.use_num_branch, -1).sum(dim=-1)
        else:
            syn_input = s_t
        delta_m_t = custom_tanh(
            self.mlp(torch.cat([syn_input, kappa_m * m_prev], dim=-1))
        m_t = kappa_m * m_prev + self.lambda_value * (1 - kappa_m) * delta_m_t
        y_t = self.w_y(m_t)
        return y_t, s_t, m_t
    def forward(self, X):
        # compute the the recurrent dynamics for a sample
        batch_size, T, num_synapse = X.shape
        kappa_m = self.kappa_m
        s_prev = torch.zeros(batch_size, len(self.tau_s), device=X.device)
        m_prev = torch.zeros(batch_size, len(self.tau_m), device=X.device)
        outputs = []
        for t in range(T):
            y_t, s_prev, m_prev = self.dynamics(X[:, t], s_prev, m_prev, kappa_m)
            outputs.append(y_t)
        return torch.stack(outputs, dim=-2)
Appendix F Results in Table Format
Table S4: The (branch) ELM on NeuronIO

### v3

@@ Appendix C Additional Results
For comparison, we fit the classic computational neuroscience models to the same NeuronIO data. Namely, we use the Leaky integrate-and-fire (LIF) neuron model (with learnable membrane timescale, weights, and bias unit for linear integration), adaptive LIF (ALIF) that has additionally a single timescale of spike-frequency adaptation. To have a fair comparison, we also fitted an ELM neuron model with only a single memory unit (and timescale) with linear synaptic integration (thus no MLP, additional parameters due to bias units, readout, and other implementation details). The LIF’s internal membrane voltage was directly fit to the target voltage, and its output spike directly to the target spikes, using otherwise same training methodology as described in Section B.
Table S4: Evaluating Tiny Models on NeuronIO
Model Soma RMSE Spike AUC # parameters
(classic) LIF 2.29 0.9117 1280
(classic) ALIF 2.29 0.9115 1281
(simplest) ELM 1.60 0.9255 1286
($d_{m}=15$) Branch-ELM 0.66 0.9915 5329
While all three models perform much worse than the reference performance threshold of $0.991AUC$, the ELM neuron performs slightly better, particularly for somatic prediction. This could result from the ELM neuron, similar to the underlying biophysical model, not enforcing an explicit hard memory reset when spiking. The noticeable lack of performance difference between LIF and ALIF might be due to the fitted models consistently staying below the spiking threshold. Finally, the LIF and ALIF model’s shortcoming in accurately capturing the I/O relationship highlights the need for a more flexible phenomenological neuron model.
@@
Figure S1: The ELM neuron is a computationally efficient model of cortical neuron. Similar figure to 2c and 2d, however, displaying FLOPs required to do inference on a single sample. a) and b) Voltage and spike prediction performance of the respective surrogate models. While previous works required around 10M parameters to make accurate spike predictions using a TCN Beniaguev et al. (2021), an LSTM baseline is able to do it with 266K, our ELM neuron model requires merely 53K, and our Branch-ELM neuron only a humble 8K, simultaneously achieving much better voltage prediction performance than the TCN. A throughput optimized ELM neuron implementation can potentially reduce the required FLOPs even further.
Figure S2: Branch-ELM neuron training is more stable with smaller $\lambda$. Evaluating a Branch-ELM with same hyper-parameters as in Figure 2, except with $\lambda=5$. The variability of test set performance is reduced for most configurations, particularly for ones with a larger number of trainable parameters (and memory units). Additionally, it allows for an improved max performance for the model type.
@@
Figure S3: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an Branch-ELM neuron with AUC $\approx 0.992$, with default hyperparameters. The number of removed divergent runs marked with $1^{*}$. a) We find above memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich (2004); Dayan & Abbott (2005). b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart & Spruston (2015); Jones & Kording (2022); Larkum (2022). c) Allowing greater updates to the memory units is beneficial, however, too large ones increase training instability. d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with the default range being 1ms-150ms. Timescales around 25ms-50ms seem to be the most useful (matching the typical membrane timescale in the cortex Dayan & Abbott (2005)); however, a lack can be partially compensated by longer timescales, even better than by the vanilla ELM. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$.
Figure S4: Ablations with the improved ELM neuron implementation. Ablations on NeuronIO that previously displayed training instabilities are now stable throughout and more consistent when rerun with the updated implementation (see section A for details on the implementation). a) Rerun of experiment in S3d. b) Rerun of experiment in 3e. c) Rerun of experiment in S3c. d) Rerun of experiment in 2c. Furthermore, we reran experiment in 3b, and training was stable for linear integration. Lastly, we reran experiment in 5e, however, did not observe significant improvements.
@@
Figure S5: The ELM neuron performs well on typical neuromorphic datasets. The following results are on the original Spiking Heidelberg Digits dataset Cramer et al. (2020). a-c) The ELM and Branch-ELM neuron reliably outperforms a classic LSTM, especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. Our LIF neuron based Spiking Neural Network (SNN), however, does manage to achieve decent performance for bin sizes around 20, in contrast to the SHD-Adding dataset (see Figure 5).
Figure S6: Ablating the ELM branch architecture on NeuronIO. Average test AUC displayed, using otherwise same hyper-parameters and training setup as in experiments in Figure 2. Exploiting the ordering in the synaptic input, and having learnable synapses is crucial for the Branch-ELM neuron model. Applying a specific nonlinearity on branch output slightly degrades performance.
Figure S7: Fitting simplified neuron models using ELM. Accurately fitting the classic Leaky Integrate and Fire (LIF) model is possible with only linear integration ($l_{mlp}=0$), and a single memory unit ($d_{m}=1$), therefore matching the ground truth architecture. When fitting both LIF and Adaptive-LIF with an ELM neuron with two memory units, the LIF fit yields better results; this is expected, as the ground truth ALIF architecture has an additional hidden state; the adaptive threshold. We suspect that as neither LIF nor ALIF display chaotic dynamics, an overall higher AUC may be achieved than for BioPhys; the AUC plateau may display residual uncertainty inherent to the dataset.
Figure S8: Ablating the $\tau_{s}$ parameter on Pathfinder-X. Average and max test Accuracy displayed, using otherwise same hyper-parameters and training setup as in experiments in Table S2. While reliably achieving high performance requires larger $\tau_{s}$, smaller timescales can achieve even higher performance, although less reliably and take longer to pick up the learning signal. The intermediate drop in performance for 75ms and 100ms could be an artifact due to nontrivial interactions between $\tau_{s}$ and the cycle length ($128$) of the flattened image ($128\times 128$) data.
Table S5: Ablation of ELM neuron $d_{m}$ on Pathfinder
$d_{m}$ 10 25 50 100 150 200 300
# params 822 3552 12k 44k 96k 168k 373k
ACC 57.83 62.95 66.31 69.31 71.54 72.98 71.85
In Table S5 we show the dependence of test accuracy on the ELM neuron model size, using otherwise same training setup as before. Performance levels out around $72\%$ accuracy at $d_{m}=150$ (default), and decreases to $57\%$ accuracy at $d_{m}=10$. An S5 model (see Table S6 with likewise a single layer (and $186K$ parameters) is outperformed with an ELM neuron with $d_{m}=25$ memory units (and $3.5K$ parameters).
Table S6: Ablation of S5 model layers on Pathfinder
# layers 1 2 4 6
# params 186k 371k 742k 1.1M
ACC 58.19 78.69 91.63 95.33
In Table S6 provide an ablation of the S5 model Smith et al. (2023), a close to state of the art model on the Long Range Arena, as reference of how such models perform with varying number of parameters and layers. The reported training hyper-parameters were used, with the learning rate individually ablated per model size. The mean accuracy over three runs is reported. Note, the steep drop-off in performance between two layers and one.
Table S7: Ablation of ELM neuron dropout probability on Pathfinder
dropout 0 0.1 0.2 0.3 0.4 0.5
Train ACC 0.8525 0.9092 0.9244 0.9149 0.8987 0.8763
Test ACC 0.7275 0.7631 0.811 0.823 0.815 0.8063
In Table S7 provide an ablation of the ELM neuron training with varying dropout probability and longer training (1000 Epochs), but otherwise same hyper-parameters as before ($d_{m}=150$). Notice, how using a dropout of 0.3 outperforms even two layered S5 at less than a third of its size. Given the significant gap between train and test accuracy, we expect that further improvements to the training setup by using weight decay, layer normalization, etc. (as routinely used in the training of SOTA models on LRA) might improve convergence speed and generalization.
@@ Appendix D Additional Visualizations
Figure S9: Visualization of ELM neuron dynamics. Extended visualization of Figure 2b for an ELM neuron achieving around 0.992 AUC. a) The synaptic input as the neuron model receives it, with excitatory input ($+1$) marked in red, and inhibitory input ($-1$) marked in blue. b) The ELM neurons predictions and the ground-truth targets for a regular sample from the data. Interestingly, the whole two seconds were inferred in one go (similar to Figure 4b), which shows its generalization capabilities beyond the training horizon of 500ms. c) A random subset of 20 synapses are visualized. Synapses receiving negative input will be deflected downwards. d) All 20 memory values are visualized. Some fluctuate more rapidly than others, typically proportional to their memory timescales.
Appendix E Results in Table Format
Table S8: The Branch-ELM on NeuronIO